OpenStudy (askme12345):

The mean score was a 75 with a standard deviation of four.....

OpenStudy (askme12345):

Can someone walk me through how to solve this 1) A statistics exam has test scores with a normal distribution. The mean score was a 75 with a standard deviation of four. -what test score gives the 95th percentile?

OpenStudy (kirbykirby):

You can standardize first. \[\frac{X-75}{4}=Z\] You can find the 95th percentile looking at a standard normal distribution table. Then you can use the above relation to see how you can get the given test score.

OpenStudy (askme12345):

okay! thanks so heres what i have. .9495 is 1.64 and .9505 is 1.65, so as far as i remember I average the 2? SO 1.645. 1.645 = x-75/4 x=81.58

OpenStudy (kirbykirby):

Yes

OpenStudy (askme12345):

Can you check this one for me? (within the same question) what is the probability that a randomly selected student has a test score between 65 and 90 z score = -2.5 for 65 and 3.75 for 90 so P = P(z=3.75) - P(z=-2.5) so .9999-.0062 gives me .9937

OpenStudy (askme12345):

Thank you sooo much, you are my statistics savior! Hahaha

OpenStudy (kirbykirby):

Yes looks good to me :)

OpenStudy (kirbykirby):

Haha I love stats